Two angles are complementary. One angle's measure is teen degrees less the three times the measure of the other. Find the measure of each of the angles using the method we used in class for word problems showing each step.
Answer by Alan3354(69443)

Two angles are complementary. One angle's measure is teen degrees less the three times the measure of the other. Find the measure of each of the angles using the method we used in class for word problems showing each step.
-----------------
I wasn't in class that day.
---------
Angles are x and y
x + y = 90 degrees
x = 3y - 10
--------------
(3y-10) + y = 90
4y - 10 = 90
4y = 100
y = 25 degs
x = 65 degs